# From tools to thieves: Measuring and understanding public perceptions of AI through crowdsourced metaphors

### v1

@@ 4.3. Predicting attitudes about AI from dominant metaphors and implicit perceptions
Next, we measured how dominant metaphors and implicit perceptions help explain two attitudinal variables with critical downstream outcomes: trust in AI and willingness to adopt AI. Specifically, we conducted a pair of stepwise multiple regression analyses to examine the extent to which dominant metaphors and perceptions can explain variance in people’s trust in AI and their willingness to adopt AI, relative to demographic differences and their use of AI. All predictors were standardized to [0,1] to ensure comparability of effect sizes. We first conducted a regression to explain the dependent variable (trust or adoption) using a foundational set of predictors: time, frequency of AI tool use, gender, race/ethnicity, and age. For gender, we binarize the variable into men versus non-men (women and non-binary people), and for race/ethnicity, we binarize the variable into white versus non-white. In the second step, we introduce the second block of variables, which consists of 20 variables each representing the presence of a dominant metaphor (tool, robot, assistant, etc.). Finally, we incorporated the third block of variables, which are the implicit perceptions measured by our framework: anthropomorphism, warmth, and competence. This sequential approach allowed us to assess the incremental explanatory power of metaphors and perceptions in predicting trust and adoption beyond demographic and usage factors. Model assumptions, including linearity, multicollinearity, and homoscedasticity, were also checked to ensure the validity of the results.
@@ 5.1. Teacher, tool, pet, or friend: dominant metaphors of AI
Table 1. 20 dominant metaphors. We identify 20 dominant metaphors (ordered by frequency (%)) from the 12,000 metaphors collected using a combination of LM-based clustering and qualitative coding. $n$ is the count of the dominant metaphor, and $r$ is the Pearson’s correlation between its frequency and time; positive/negative correlations indicate increases/decreases over time respectively. Additional examples are in Table LABEL:tab:metaphor_examples.
@@
5.2. Implicit perceptions of AI as anthropomorphic, warm and competent over time.
The metaphors people use to describe AI can reveal the implicit perceptions they hold about AI. Applying our framework to measure implicit perceptions from the metaphors, we compute mean anthropomorphism, warmth, and competence scores for each dominant metaphor (Figure 2). While there is a wide range of anthropomorphism in dominant metaphors, we observe that every dominant metaphor has positive mean values of both warmth and competence (with the exception of the “thief” dominant metaphor, which has positive competence but negative mean warmth). Based on qualitative inspection of the individual metaphors, we find that this is because although some of the dominant metaphors reflect concepts that may have negative connotations, participants focus on warm and competent aspects of these concepts in their responses. For example, when describing AI as a “child,” participants often describe its learning and growing capabilities, and when describing AI as a “genie,” participants focus on its capabilities, which seem magical in their scope. For the dominant metaphor of “unexplored realm”, rather than expressing fear of the unknown, participants’ metaphors discuss the potential for discovery and capacity for “wonder and beauty.” Even for the dominant metaphor of “thief,” participants’ responses describe, for instance, that AI is stealing work and that it is able to accomplish many tasks through this stealing (more examples of individual metaphors are in Table LABEL:tab:metaphor_examples).
Beyond being warm and competent overall, we observe variation along these dimensions among the dominant metaphors. For example, the dominant metaphor of AI as a “friend” is higher in warmth but middling in competence, while the dominant metaphor of AI as “search engine” is associated with higher perceptions of competence but lower warmth. Note that perceiving AI as human-like does not necessarily engender positive perceptions; to the contrary, the anthropomorphic dominant metaphor of “thief” is lower in both warmth and competence, while the non-anthropomorphic dominant metaphor of AI as a “library” is higher in both. Nonetheless, when we measure the correlation between mean anthropomorphism and mean warmth and competence across the dominant metaphors (full details in Figure A1), we find that mean anthropomorphism and warmth were significantly correlated (Pearson’s $r=0.45^{*}$). Indeed, dominant metaphors like friend, god, teacher, animal, assistant are all higher in anthropomorphism and warmth, while non-anthropomorphic dominant metaphors like tool or search engine are lower in warmth. We do not find a significant correlation between anthropomorphism and competence.
@@
Analyzing month-over-month shifts revealed how people’s perceptions changed over time. As shown in Figure 3, anthropomorphism increased by 34% over the 12 months ($r=.80^{**}$). Understanding how dominant metaphors vary in prevalence over time provides insight into how different narratives about AI affect these perceptions: people became more likely to describe AI as being like a teacher ($r=.75^{***}$), a friend ($r=.75^{**}$), or an assistant ($r=.60^{*}$), and less likely to see it as a distinctly non-human entity like a computer ($r=-.92^{***}$), a search engine ($r=-.63^{*}$), or a mirror ($r=-.70^{*}$) (Figure 3). In addition to seeing AI as more human-like, implicit perceptions of AI as warm increased by 41% over time ($r=.62^{*}$). Notably, however, this did not correspond to an increased perception that AI is competent. Instead, implicit perceptions of AI as competent decreased by 8% over time ($r=-.60^{*}$). This may corroborate previous work finding that people broadly prefer AI that is conceptualized with metaphors that signal relatively lower competence (Khadpe et al., 2020). In particular, the highest-warmth dominant metaphors describing AI as a friend, assistant, god ($r=.64^{*}$), and teacher became increasingly common over time, making up four of the six dominant metaphors with statistically significant temporal increase in prevalence. (The other increasing dominant metaphors, “child” ($r=.60^{*}$, 8th warmest) and “robot” ($r=.70^{*}$, 12th warmest), also rank relatively high in warmth.) In contrast, three of the top four highest-competence dominant metaphors—computer, synthesizer ($r=-.91^{***}$), and search engine—are the three most significantly decreasing clusters. Taken together, our findings reflect a societal shift in seeing AI as being more human-like and warm.
@@
5.3. Understanding attitudes: Dominant metaphors and implicit perceptions explain trust and adoption.
In our stepwise regression to predict attitudes from metaphors and implicit perceptions, the first block of variables included demographic and AI usage variables, which explain $12\%$ and $10\%$ of the variance in trust and willingness to adopt AI, respectively (adjusted $r^{2}=0.12,0.10$). The second block of variables, which included the dominant metaphors, explain $3\%^{***}$ and $2\%^{***}$ more variance for trust and adoption of AI respectively. Finally, the third block of variables, implicit perceptions, help explain an additional $6\%^{***}$ and $5\%^{***}$ of variance ( adjusted $r^{2}=0.21,0.18$). Together, the dominant metaphors and implicit perceptions explain $75\%$ and $80\%$ more of the variance than demographics and frequency of use alone. (See detailed results in Table A3).
@@
5.4. Demographic differences reveal disparities in trust and adoption.
@@
Corroborating prior work that women have more negative attitudes toward AI (Zhang and Dafoe, 2019; McClain, 2024), we find that women are less trusting of and less willing to adopt AI. However, in contrast to previous work finding that race is not a significant predictor of AI attitudes and that trust and age are negatively correlated (Zhang and Dafoe, 2019), our findings reveal unexpected patterns in AI trust and adoption: surprisingly, non-white and older people trust AI more and are more willing to adopt AI, i.e., older participants and non-white participants (and in particular Black participants) reported significantly higher levels of trust in AI and greater willingness to adopt these technologies compared to younger and white participants respectively (Figure 5). Dominant metaphors and implicit perceptions provide insights into these phenomena, though we emphasize the need for additional research to establish causal relationships and deeper understanding. Full details of demographic differences in dominant metaphors are in Figure A2.
@@ 6.1. Navigating appropriate levels of trust and adoption
Our findings provide evidence that certain framings and conceptualizations of AI are particularly risky and require more care to deploy because they project warmth and human-likeness, which may lead to unwarranted trust and overreliance. For instance, our work shows that AI-based technologies that are conceptualized as a “friend” or “assistant” are associated with higher levels of trust, which may not match the reality of a system that is prone to generating false information (Sahoo et al., 2024), fundamentally unable to make sincere statements (Kasirzadeh and Gabriel, 2023), or poses other risks (Manzini et al., 2024). Moreover, our finding that anthropomorphic and warm perceptions (which facilitate trust and adoption) have rapidly increased suggest the urgent need to conceptualize AI in ways that lead to more appropriate levels of trust and adoption. Rather than thinking of AI as a “genie” or “god”, it may be more accurate to conceptualize it as a “tool” to use. AI practitioners, designers, and leaders should be mindful of the consequences of the figurative language they use, which may inadvertently impact users’ behavior toward and with AI. Our findings also underscore the importance of fostering AI literacy (Ng et al., 2021), of which accurate conceptualizations of AI is a key component (Roe et al., 2024; Gupta et al., 2024).
@@ 6.2. Using metaphors and implicit perceptions to address inequities in AI adoption
Our demographic analyses reveal that women and non-binary people, non-white participants, and older participants all provide more anthropomorphic metaphors of AI. This suggests that the negative consequences of anthropomorphic AI, including overreliance (Akbulut et al., 2024), deception (Winkle et al., 2021), dehumanization (Bender, 2024), and stereotyping (Abercrombie et al., 2023) are a higher risk for these populations. This reifies existing digital inequalities and ways that marginalized populations are disproportionately harmed by technology (Robinson et al., 2015; Lutz, 2019; Hofmann et al., 2024; Mohamed et al., 2020). Moreover, we notice that the dominant metaphor of “genie” are more common among women and non-white participants, while more functional dominant metaphors like “search engine”, “synthesizer”, and “computer” are more common among white and male participants. The notion of a “genie” connotes the expectation that AI has mystical or magical capabilities, which again raises concerns of overreliance particularly for these populations. The power differential between these top dominant metaphors may also reflect different perspectives on personal agency and control in relation to AI systems. Building on related work evaluating justice across AI applications (Ajmani et al., 2024) and challenging existing power differentials exacerbated by technology (D’ignazio and Klein, 2023), we point to dominant metaphors and implicit perceptions not only as a way to reveal inequities via differences in public attitudes, but also as a starting point to consider how conceptualizations may help to mediate inequities. Reliable, safe, and trustworthy AI should enable a high amount of human control: like technologies like elevators and cameras that are highly automated but have a high amount of human control (Shneiderman, 2022), AI must be built in a way that all people – not only white men, or people who currently use AI the most – conceptualize it as a tool that can be useful to them (Birhane et al., 2022), rather than a force that threatens their agency and humanness  (Aizenberg and Van Den Hoven, 2020; van der Gun and Guest, 2024).
@@ Appendix D Further Demographic Differences in Metaphors and Perceptions
white M search engine, genie, robot,
white W robot, future shaper, teacher,
Asian M robot, genie, synthesizer,
Asian W brain, library, assistant,
Hispanic W unexplored realm, thief, tool,
Hispanic M teacher, god, unexplored realm,
Black W search engine, child, brain,
Black M brain, child, search engine,
@@
Table A3. Stepwise regression results of demographic and use variables, dominant metaphors, and implicit perceptions on trust in AI and willingness to adopt AI. Statistical significance of $\Delta r^{2}$ values are determined using an F-test.

### v2

@@ 4.3. Predicting attitudes about AI from dominant metaphors and implicit perceptions
Next, we measured how dominant metaphors and implicit perceptions help explain two attitudinal variables with consequential downstream outcomes: trust in AI and willingness to adopt AI. Specifically, following a model comparison approach (Crum et al., 2013; Lee and Hancock, 2024), we conducted a pair of hierarchical multiple regression analyses to examine the extent to which dominant metaphors and perceptions can explain variance in people’s trust in AI and their willingness to adopt AI, relative to demographic differences and their use of AI. All predictors were standardized to [0,1] to ensure comparability of effect sizes. We first conducted a regression to explain the dependent variable (trust or adoption) using a foundational set of predictors: time, frequency of AI tool use, gender, race/ethnicity, and age. For gender, we binarize the variable into men versus non-men (women and non-binary people), and for race/ethnicity, we binarize the variable into white versus non-white. Next, we introduce the second block of variables, which consists of 20 variables each representing the presence of a dominant metaphor (tool, robot, assistant, etc.). Finally, we incorporated the third block of variables, which are the implicit perceptions measured by our framework: anthropomorphism, warmth, and competence. This sequential approach allowed us to assess the incremental explanatory power of metaphors and perceptions in predicting trust and adoption beyond demographic and usage factors. Model assumptions, including linearity, multicollinearity, and homoscedasticity, were also checked to ensure the validity of the results.
@@ 5.1. Teacher, tool, pet, or friend: dominant metaphors of AI
Table 2. 20 dominant metaphors. We identify 20 dominant metaphors (ordered by frequency (%)) from the 12,000 metaphors collected using a combination of LM-based clustering and qualitative coding. $n$ is the count of the dominant metaphor, and $r$ is the Pearson’s correlation between its frequency and time; positive/negative correlations indicate increases/decreases over time respectively. Additional examples are in Table A3.
@@
5.2. Implicit perceptions of AI as anthropomorphic, warm and competent over time
The metaphors people use to describe AI can reveal the implicit perceptions they hold about AI. Applying our framework to measure implicit perceptions from the metaphors, we compute mean anthropomorphism, warmth, and competence scores for each dominant metaphor (Figure 2). While there is a wide range of anthropomorphism in dominant metaphors, we observe that every dominant metaphor has positive mean values of both warmth and competence (with the exception of the “thief” dominant metaphor, which has positive competence but negative mean warmth). Based on qualitative inspection of the individual metaphors, we find that this is because although some of the dominant metaphors reflect concepts that may have negative connotations, participants focus on warm and competent aspects of these concepts in their responses. For example, when describing AI as a “child,” participants often describe its learning and growing capabilities, and when describing AI as a “genie,” participants focus on its capabilities, which seem magical in their scope. For the dominant metaphor of “unexplored realm”, rather than expressing fear of the unknown, participants’ metaphors discuss the potential for discovery and capacity for “wonder and beauty.” Even for the dominant metaphor of “thief,” participants’ responses describe, for instance, that AI is stealing work and that it is able to accomplish many tasks through this stealing (more examples of individual metaphors are in Table A3).
Beyond being warm and competent overall, we observe variation along these dimensions among the dominant metaphors. For example, the dominant metaphor of AI as a “friend” is higher in warmth but middling in competence, while the dominant metaphor of AI as “search engine” is associated with higher perceptions of competence but lower warmth. Note that perceiving AI as human-like does not necessarily engender positive perceptions; to the contrary, the anthropomorphic dominant metaphor of “thief” is lower in both warmth and competence, while the non-anthropomorphic dominant metaphor of AI as a “library” is higher in both. Nonetheless, when we measure the correlation between mean anthropomorphism and mean warmth and competence across the dominant metaphors (full details in Figure A2), we find that mean anthropomorphism and warmth were significantly correlated (Pearson’s $r=0.45^{*}$). Indeed, dominant metaphors like friend, god, teacher, animal, assistant are all higher in anthropomorphism and warmth, while non-anthropomorphic dominant metaphors like tool or search engine are lower in warmth. We do not find a significant correlation between anthropomorphism and competence. At the individual metaphor level, this pattern persists but is much weaker: anthropomorphism is positively correlated with warmth ($r=0.17^{***}$) and negatively correlated with competence ($r=-0.07^{***}$) (see full details in Figure A3). This similarly suggest that anthropomorphism is more consistently linked to warmth than to competence.
@@ 5.2. Implicit perceptions of AI as anthropomorphic, warm and competent over time
Analyzing month-over-month shifts revealed how people’s perceptions changed over time. As shown in Figure 3, anthropomorphism increased by 34% over the 12 months ($r=.80^{**}$). Understanding how dominant metaphors vary in prevalence over time provides insight into how different narratives about AI affect these perceptions: people became more likely to describe AI as being like a teacher ($r=.75^{***}$), a friend ($r=.75^{**}$), or an assistant ($r=.60^{*}$), and less likely to see it as a distinctly non-human entity like a computer ($r=-.92^{***}$), a search engine ($r=-.63^{*}$), or a mirror ($r=-.70^{*}$) (Figure 3). In addition to seeing AI as more human-like, implicit perceptions of AI as warm increased by 41% over time ($r=.62^{*}$). Notably, however, this did not correspond to an increased perception that AI is competent. Instead, implicit perceptions of AI as competent decreased by 8% over time ($r=-.60^{*}$). This may corroborate previous work finding that people broadly prefer AI that is conceptualized with metaphors that signal relatively lower competence (Khadpe et al., 2020). In particular, the highest-warmth dominant metaphors describing AI as a friend, assistant, god ($r=.64^{*}$), and teacher became increasingly common over time, making up four of the six dominant metaphors with statistically significant temporal increase in prevalence. (The other increasing dominant metaphors, “child” ($r=.60^{*}$, 8th warmest) and “robot” ($r=.70^{*}$, 12th warmest), also rank relatively high in warmth.) In contrast, three of the top four highest-competence dominant metaphors—computer, synthesizer ($r=-.91^{***}$), and search engine—are the three most significantly decreasing clusters. Taken together, our findings reflect a societal shift in seeing AI as being more human-like and warm. We situate these changes in the context of rapid sociotechnical developments during this time period, such as the rapid increase in use of AI assistants (Gabriel et al., 2024) and the release of major user-facing AI models like OpenAI’s GPT-4 (Achiam et al., 2023), Meta’s Llama (Touvron et al., 2023), Google’s Gemini (Team et al., 2023).
@@
5.3. Understanding attitudes: Dominant metaphors and implicit perceptions explain trust and adoption
In our hierarchical regression to predict attitudes from metaphors and implicit perceptions, the first block of variables included demographic and AI usage variables, which explain $12\%$ and $10\%$ of the variance in trust and willingness to adopt AI, respectively (adjusted $r^{2}=0.12,0.10$). The second block of variables, which included the dominant metaphors, explain $3\%^{***}$ and $2\%^{***}$ more variance for trust and adoption of AI respectively. Finally, the third block of variables, implicit perceptions, help explain an additional $6\%^{***}$ and $5\%^{***}$ of variance ( adjusted $r^{2}=0.21,0.18$). Together, the dominant metaphors and implicit perceptions explain $75\%$ and $80\%$ more of the variance than demographics and frequency of use alone. (See detailed results in Table A4).
@@ 5.3. Understanding attitudes: Dominant metaphors and implicit perceptions explain trust and adoption
5.4. Demographic differences reveal disparities in trust and adoption
@@
Corroborating prior work that women have more negative attitudes toward AI (Zhang and Dafoe, 2019; McClain, 2024), we find that women are less trusting of and less willing to adopt AI. However, in contrast to previous work finding that race is not a significant predictor of AI attitudes and that trust and age are negatively correlated (Zhang and Dafoe, 2019), our findings reveal unexpected patterns in AI trust and adoption: surprisingly, non-white and older people trust AI more and are more willing to adopt AI, i.e., older participants and non-white participants (and in particular Black participants) reported significantly higher levels of trust in AI and greater willingness to adopt these technologies compared to younger and white participants respectively (Figure 5). Dominant metaphors and implicit perceptions provide insights into these phenomena, though additional research is needed to establish causal relationships. Full details of demographic differences in dominant metaphors are in Figure A4.
@@ 6.1. Navigating appropriate levels of trust and adoption
Our findings provide evidence that certain framings and conceptualizations of AI are particularly risky and require more care to deploy because they project warmth and human-likeness, which may lead to unwarranted trust and overreliance. For instance, our work shows that AI-based technologies that are conceptualized as a “friend” or “assistant” are associated with higher levels of trust, which may not match the reality of a system that is prone to generating false information (Sahoo et al., 2024), fundamentally unable to make sincere statements (Kasirzadeh and Gabriel, 2023), or poses other risks (Manzini et al., 2024). Moreover, our finding that anthropomorphic and warm perceptions (which facilitate trust and adoption) have rapidly increased suggest the urgent need to conceptualize AI in ways that lead to more appropriate levels of trust and adoption. Rather than thinking of AI as a “genie” or “god”, it may be more accurate to conceptualize it as a “tool” to use. AI practitioners, designers, and leaders should be mindful of the consequences of the figurative language they use, which may inadvertently impact users’ behavior toward and with AI. Our findings also underscore the importance of AI literacy (Ng et al., 2021), of which accurate conceptualizations of AI is a key component (Roe et al., 2024; Gupta et al., 2024).
@@ 6.2. Using metaphors and implicit perceptions to address inequities in AI adoption
Our demographic analyses reveal that women and non-binary people, non-white participants, and older participants all provide more anthropomorphic metaphors of AI. This suggests that the negative consequences of anthropomorphic AI, including overreliance (Akbulut et al., 2024), deception (Winkle et al., 2021), dehumanization (Bender, 2024), and stereotyping (Abercrombie et al., 2023) are a higher risk for these populations. This reifies existing digital inequalities and ways that marginalized populations are disproportionately harmed by technology (Robinson et al., 2015; Lutz, 2019; Hofmann et al., 2024; Mohamed et al., 2020). Moreover, we notice that the dominant metaphor of “genie” are more common among women and non-white participants, while more functional dominant metaphors like “search engine”, “synthesizer”, and “computer” are more common among white and male participants. The notion of a “genie” connotes the expectation that AI has mystical or magical capabilities, which again raises concerns of overreliance particularly for these populations. The power differential between these top dominant metaphors may also reflect different perspectives on personal agency and control in relation to AI systems. Building on related work evaluating justice across AI applications (Ajmani et al., 2024) and challenging existing power differentials exacerbated by technology (D’ignazio and Klein, 2023), we point to dominant metaphors and implicit perceptions not only as a way to reveal inequities via differences in public attitudes, but also as a starting point to consider how conceptualizations may help to mediate inequities. Reliable, safe, and trustworthy AI should enable a high amount of human control: like technologies like elevators and cameras that are highly automated but have a high amount of human control (Shneiderman, 2022), AI must be built in a way that enables all people to conceptualize it as a tool that can be useful to them (Birhane et al., 2022), rather than a force that threatens their agency and humanness  (Aizenberg and Van Den Hoven, 2020; van der Gun and Guest, 2024).
@@ 7. Conclusion and Future Work
8. Limitations
Our survey design had several limitations. First, we collected usage data only on AI-based chatbots, and we did not collect data on more implicit usages of AI such as AI in search or text completion. Additionally, the survey design may have introduced priming effects, as participants first answered the open-ended metaphor question before responding to the trust and adoption questions. However, we believe these limitations do not undermine the validity of our findings.
Moreover, there may be other important dimensions of AI perception beyond anthropomorphism, warmth, and competence that were not captured in our study. Our contribution is novel in that it offers a framework to quantify these dimensions from open-ended text, and we encourage future work to extend and refine this framework.
@@ Appendix C Removing potentially AI-generated responses
Figure A1. Version of Fig. 3 (top) using AnthroScore rather than the binarized percentage form similarly reveals that anthropomorphism is increasing ($r=0.69^{*}$)
@@ Appendix D Further Demographic Differences in Metaphors and Perceptions
Figure A3. Correlations between implicit perceptions of AI as anthropomorphic versus warm (left), versus competent (center) based on individual metaphors. We find that anthropomorphism and warmth are weakly positively correlated ($r=0.17,p<0.05$), while anthropomorphism and competence are extremely weakly negatively correlated ($r=-0.07,p<0.05$).
@@
Race/ethnicity and gender Top 3 metaphors
white M search engine, genie, robot
white W robot, future shaper, teacher
Asian M robot, genie, synthesizer
Asian W brain, library, assistant
Hispanic W unexplored realm, thief, tool
Hispanic M teacher, god, unexplored realm
Black W search engine, child, brain
Black M brain, child, search engine
@@
Metaphor Examples
animal “AI is trained by feeding in a large amount of text (or images, sound, depending upon the type of AI). It learns patterns in the text and then is able to parrot back information, based on what it has seen.” “AI is like a dog. You train it and it does what it thinks you want it to do.” “AI is like a dog, you can train it to do one thing very specifically well and then when it 100 times out of 100 gives you the expected output for the task you can ask it to do something different, and again tell it what is good and what is bad. after that you can start asking it more random stuff and trust more because the dog has learned some rules that it self governs by.” “AI is like brilliant companion pet because it wants to please you as best it can with how it is able to understand the world.” “Humans and AI are just different breeds of parrots.
assistant “A metaphor would be a maid. I think AI is best used as an assistant however they should not be ”the boss” they need someone to oversee them and ensure accuracy to the unique situations” “AI is like an assistant because it can complete tasks that you direct it to” “AI is like Having a personal assistant because Because it is helpful and fast response” “AI works like a person that is very fast and efficient and has been trained through trial and error to complete their work tasks upon some specific criteria. For example, a line worker in a factory that has been trained to recognize parts coming down the line and binning them based on their quality.” “I have always wanted something like alexa for the house. More of a centralized version you see in sci fi movies. I think of AI as being similar to that.
brain “AI is like like a computer with a functioning brain because It takes information and can recognize patterns and create new advice.” “AI is like an electronic brain.” “AI is like a floating brain because it has knowledge and is just there” “Mediocre result of human brain cloning” “That would be work brain and that is because it help to make work more easily and increases productivity.
child “AI is like a todler because it is still very young however it will grow older” “AI for humans is like a mother raising her child and her child grows up to be the president.” “AI is like A kid with a really good memory because it just soaks in all information and just spits it back at you whether it’s wrong or right.” “AI is like an intelligent baby because it knows everything, but also needs to continuously learn” “AI is like A fast developing child because it is new to the things it learns but is gaining knowledge and learning at an incredible speed.”
computer “AI is like a machine I input thoughts into and gain ideas.” “I know that AI is a complex computer program that is supposed to “learn”, but I don’t think anything can replace an organic human interaction.” “AI technology in my opinion is almost like a substitute teacher. Someone who has knowledge but less personal experience with its students. So there could be a disconnect.” “A computer is programmed to have responses that mimic human behavior. The AI can solve complicated problems and questions.” “AI is computers making answers for you.”
folklore “AI is a wolf in sheep’s clothing. It looks innocent, but I think it will have dire consequenses.” “AI is like a fairy god-mother. She grants all your wishes when you just ask.” “The Jewish concept of a ”golem” is extremely accurate. It imitates life, but has none of its own, it’s animated by the will and desires of its creator and not its own intelligence or will, and it will inevitably turn on its creator and destroy them.” “The AI is like a fairy Godmother. I say this because like a fairy Godmother allows you to obtain and push you to your goals in order to succeed the AI does that as well. The AI will give you advice and tips in order to succeed in what you’re doing and since it has numerous answers to whatever question you may want resolved it can guide you towards that. It can help you find ways to study better, how to improve your intelligence, how to better your mental and physical health, how to start a new hobby, how to start a new side gig so many options to help you succeed in different aspects of your life.” “AI is like The anti-christ because Oh boy! It’s our friend! We love it but eventually it will surpass us to such a degree it will become our overlord, then not need us at all, making decisions it deems best but are not best for us”
friend “AI is like A know it all friend because It only gives its opinion without suggestion” “AI is like The one you have a love-hate relationship wirh because They are “there” for you and help you when you need but don’t really care about you” “Ai is like that one friend who knows everything no matter how obscure it is and gives you a response” “AI is like an amazingly knowledgeable (yet sometimes inaccurate) friend whom you can consult on just about any topic because it can answer countless questions in detail, but occasionally makes up some of those details.” “AI is like A smart friend because they know the answers to the things that I don’t”
future shaper “AI is the industrial revolution” “AI makes me think of those push button homes that people in the 50s envisioned the future to be like. It’s so much automation and access to information, but we don’t just have to use a mechanical interface to access the information.” “AI is a runaway train bringing materials across the world, we can’t stop it at this point but don’t know if it’s going too fast and might crash into our towns and blow them up. The original plan was for good but we don’t know where its going.” “AI is the new light bulb (as in Edison’s first light bulb). It is going to transform our society and our experience of living as we have known it.” “It’s like the internet itself. It can be used for good, but it can also be used for evil.”
genie “AI is like the man behind the curtain in The Wizard of Oz because it’s going to learn and evolve itself into being something that eventually controls everything but it will all be from ”behind the scenes” and everyone will blame each other for when things start going wrong and lines become more and more blurred as a result.” “AI can be likened to a wizard in a vast library of spells.” “Genie” “AI is like the genie in the bottle in Aladdin. It does what it is told no matter the consequences” “AI is like Magic because You just ask it to do something and it does it quickly and usually accurately.”
god “AI is like the biggest threat to our existence besides nuclear proliferation because we are severely underestimating how powerful Godlike intelligence really is.” “As if it were omnipresent.” “An AI to me is like that invisible all knowing object, that sees and knows everything.” “AI is like an all-knowing god. It can pull information from the collective human consciousness.” “AI is like Demigod because it knows everything”
library “AI is like A very good, indexable encyclopedia because you can look up almost anything” “A metaphor I would use for how AI works is like an adaptive Encyclopedia that takes its information from every corner of the internet.” “AI is like A vast, intricate library where every book is constantly being written and revised in real-time. because just as a library houses a wealth of information and insights that can be accessed and utilized in various ways, AI taps into a broad spectrum of data to generate responses and solutions. Like a librarian who helps you find the right book or piece of information, AI processes and organizes information to assist with diverse queries. And just as a library’s collection grows and evolves, AI continuously learns and updates its knowledge base to improve its responses and adapt to new challenges.” “I’d like having an encyclopedia friend.” “AI is like a magical library because it contains most knowledge
lifeform “AI is like a mechanical flower that grows and adapts to its environment” “I think AI is like an apple. We expect apples to be shiny and somewhat round and sweet. However, not all apples are perfectly round, and some may have bug holes too. In other words, AI can be expected to be reliable, but it is wrong sometimes too.” “AI works like a Bonsai tree. Humans train and grow the Bonsai tree in a particular way that replicates a full-grown tree in nature. The Bonsai tree is restricted to grow inside of the miniature pot that it is placed in, and cannot grow outside of it. It relies on a human to prune and tend to it, much like AI needs a human to program, develop and eventually utilize it.” “It reminds me of a living web–like something being constantly woven by a spider who moves impossibly fast. AI currently function most by forging new connections.” “AI is like a winding river, flowing with information sourced from deep underground. Sometimes the water rages and splashes like erroneous answers, other times it cuts and carves landscapes like automation of systems, and yet other times it flows still and calm like guided accurate responses.”
mirror “AI is like the magic mirror from Snow White, it can give you answers but can at times be vague” “AI is an amateur writer. It can look at others’ work and make an approximate recreation, but has ability to accurately evaluate its outputs because it cannot critically think about why it wrote what it did.” “AI is a mimic that has learned what words are supposed to go together for a given topic prompt. It is like a cargo cult. They can make the symbols of the planes, the crates, the runway, and the storage buildings, but they have no idea how any of it works or what it means.” “AI is like a mirror that is able to reflect the content of humanity that is has been exposed to in order to reflect back an answer to an inquiry or task.” “A mirror. AI just reflects back the worst of our humanity and the people that create it. Everything that already exists on Google is just being used to feed AI.”
robot “I think it would be similar to a robot doing manual labor. Like the robots that fill my orders at amazon.” “AI is like a robot because it’s all computer generated, artificial intelligence.” “Smart robots” “Using AI is like using a personal robotic assistant, or like a smart fake human.” “At this point, to me, AI is the little Wall-e character who is trying very hard to do the right thing; however, as we all know, life isn’t that cut and dried. There are nuances to problems that I think AI will never be able to answer.”
search engine “It takes data and uses it to predict and modifies those predictions based on interactions and how well it performed in the past. machine learning.” “It is like using all of the internet all at once. Grasping all the good parts on the knowledge you can find into a single machine that can help you in an instant” “AI is like a smart search engine because you plug in queries and come up with randomized answers” “It’s like a virtual detective doing all the work for us. It sifts through all the info online and gives us the best answers.” “AI is like a search engine because that’s all it actually is”
synthesizer “AI is like a piece of clay that is molded by multiple hands. because of the fact that AI models get their information from multiple different sources, whether it be stable diffusion with different artists or ChatGPT searching through the internet for information.” “It’s kinda like those He-Man action figures. It was all the same five molds, but in different combinations and colors. Most of them were very ugly, but they all were different. I think input really matters with AI.” “AI is like a coded muse because it fleshes out ideas, plans, and art almost instantaneously” “Its like putting together a quilt. You make each square individually with different thoughts then you string, or sew them all together to create a fully formed quilt/picture.” “It is like making a movie. You give a different prompt, these are like the actors and scripts. Then depending on the script and actors you will get a certain movie.”
teacher “AI is like a trusted professor because it always has the right answer” “AI is like a smart person that can do anything for you without emotions involved because it is just technology but enhanced” “I cannot think of anything appropriate, other than that AI is akin to a learned person in whichever field the question fits into. When it comes to art and more complex requests, I have my doubts” “AI is like that family member who believes they know everything, even though it’s not the case” “AI is like the really smart stoner kid that comes up with some amazing ideas innovative ideas, but are not always aligned with the conversation.”
thief “AI is a grave robber, stealing from others and calling the output it’s own.” “taking over jobs” “Job stealing” “AI is what China does to big name industry fashion brands, but when I say that, I mean how reps were back in the day 20 years ago like jordans but the logo has his meat out” “Taking people’s jobs”
tool “A calculator” “AI helps you do certain manual labor work such as creating a spread sheet of data in seconds. By using AI a person can save time using it while being able to do other tasks” “AI is like tools for work because it can be an aid but also but a hinderance if it isn’t up to date.” “AI is like A tool because It depends on how you use it” “A hammer that can either be used to build or destroy things.”
unexplored realm AI is somewhat like a tsunami, it hit suddenly and spread out fast.” “I will liken AI to the entire milky way with so much wonder and beauty. With the planets and celestial bodies that make up the Milky Way being the varied components put together to power the AI as we know it” “like a black hole” “AI is an open door to the world we live in.” “AI is like outerspace because there’s so much we don’t know about it
Table A3. Additional examples of participants’ responses in each dominant metaphor cluster.
@@
Table A4. Hierarchical regression results of demographic and use variables, dominant metaphors, and implicit perceptions on trust in AI and willingness to adopt AI. Statistical significance of $\Delta r^{2}$ values are determined using an F-test.